# Fluid Antenna-enabled Near-Field Integrated Sensing, Computing and Semantic Communication for Emerging Applications

### v1

@@ II-A Signal Transmission Model
The semantic communication flow in the proposed framework is shown in Fig. 3. In semantic communication, semantic extraction is primarily guided by the local KBs of the sender and receiver. These local KBs are constructed from a shared KB and enriched with individual knowledge unique to each party. For example, in a vehicular network, the shared KB could be a global map, while each vehicle’s local KB includes this global map along with specific local routes the vehicle has previously traversed. Based on the local KB, the sender focuses on extracting and conveying the underlying meaning of a message. To achieve this, a machine learning-based semantic extraction module is employed to identify key elements that capture the core semantic features of the message. For text-based messages, this typically involves leveraging models such as Transformers to select keywords that maximise performance metrics, such as the similarity between the original message and the reconstructed message at the receiver. The selected words are then encoded using channel coding techniques to ensure robustness against channel noise and transmission errors. A more integrated approach in semantic communication is to employ joint source-channel coding (JSCC), which is a machine learning-based approach such as proposed in [28, 39], where the transmitted signal $\mathbf{C}$ is generated through a JSCC module $\mathcal{F}(\mathbf{S})$, with $\mathbf{S}$ denoting the source message. Therefore, for each user $k$, the semantic compression ratio is given by
@@
At the receiver’s end, the local KB enables accurate reconstruction of the sender’s intended meaning, even if the exact transmitted words are not received. It is important to note that semantic noise plays a critical role in influencing the performance of semantic inference. Semantic noise may arise from poor channel conditions that cannot be fully mitigated by channel coding and decoding techniques, resulting in semantic distortion. It may also stem from impaired or incomplete KBs, which can lead to misinterpretation of the message [40]. To address the former, the proposed framework employs FAs to enhance channel conditions and reduce the impact of transmission-related semantic noise. Semantic noise caused by impaired KBs is beyond the scope of this work, and we assume that the KBs are properly constructed and maintained.
@@
where $\mathbf{W}_{c}\in\mathbb{C}^{(N_{\text{tx}}\times N_{\text{tz}})\times K}$ denote the communication beamforming matrix for the CUs, and each column corresponds to the beamforming vector for an individual user. Specifically, $\mathbf{w}_{k}\in\mathbb{C}^{(N_{\text{tx}}\times N_{\text{tz}})\times 1}$ represents the beamforming vector for the $k$-th user. Thus, the complete beamforming matrix can be expressed as $\mathbf{W}_{c}=[\mathbf{w}_{0},\mathbf{w}_{1},\cdots,\mathbf{w}_{K-1}]$. The semantic communication signals intended for the CUs are denoted by $\mathbf{C}\in\mathbb{C}^{F\times K}$, where each column represents the signal intended for a specific user across $F$ frames. Specifically, $\mathbf{c}_{k}\in\mathbb{C}^{F\times 1}$ denotes the communication signal for the $k$-th CU. Accordingly, the communication signal matrix is given by $\mathbf{C}=[\mathbf{c}_{0},\mathbf{c}_{1},\cdots,\mathbf{c}_{K-1}]$. Additionally, $\mathbf{W}_{r}\in\mathbb{C}^{(N_{\text{tx}}\times N_{\text{tz}})\times F}$ represents the sensing signal. Without loss of generality, if the number of frames $F$ is sufficiently large, we can make the following two assumptions [41]:
@@ II-B Communication Model
where $\mathbf{\Phi}_{k}\in\mathbb{C}^{1\times(N_{\text{tx}}\times N_{\text{tz}})}$ represents the free space path-loss vector for the $k$-th user, where each element is give by $\frac{1}{\sqrt{4\pi||\mathbf{u}_{i}-\mathbf{b}_{k}||^{2}}}$ under a non-uniform spherical wave (NUSW) model [7], and $\mathbf{a}_{t}\left(\theta_{k},\phi_{k},d_{k},\mathbf{u}\right)\in\mathbb{C}^{1\times(N_{\text{tx}}\times N_{\text{tz}})}$ denotes the steering vector. Additionally, $\mathbf{n}_{k}\sim\mathcal{CN}(0,\sigma^{2}_{c}\mathbf{I}_{1\times F})$ represents the communication noise.
@@
where $N_{s}$ is the number of scatterers forming an extended target, and the subscript $s$ indicates the $s$-th scatterer. Additionally, $\mathbf{\Phi}_{l,s}\in\mathbb{C}^{1\times(N_{\text{tx}}\times N_{\text{tz}})}$ is the path-loss vector where each element is given by $\frac{1}{\sqrt{4\pi||\mathbf{u}_{i}-\mathbf{b}_{l,s}||^{2}}}$, and $\mathbf{a}_{t}\left(\theta_{l,s},\phi_{l,s},d_{l,s},\mathbf{u}\right)\in\mathbb{C}^{1\times(N_{\text{tx}}\times N_{\text{tz}})}$ denotes the corresponding steering vector for the target. The term $\mathbf{n}_{l}\sim\mathcal{CN}\left(0,\sigma^{2}_{c}\mathbf{I}_{1\times F}\right)$ represents the noise encountered by the target. The received signal by target $l$ related to user $k$ is given by:
@@
Given that both the CUs and targets are located within the NF region of the BS, the steering vector can be characterised by [7]:
@@ II-C Sensing Model
where $\mathbf{N}\sim\mathcal{CN}\left(0,\sigma^{2}_{r}\mathbf{I}_{(N_{\text{rx}}\times N_{\text{rz}})\times F}\right)$ represents the Gaussian noise, $\mathbf{\Psi}_{l,s}\in\mathbb{C}^{(N_{\text{rx}}\times N_{\text{rz}})\times(N_{\text{tx}}\times N_{\text{tz}})}$ represents the path loss matrix with each element being $\frac{1}{4\pi||\mathbf{u}_{i}-\mathbf{b}_{l,s}||||\mathbf{v}_{m}-\mathbf{b}_{l,s}||}$, and $\mathbf{a}_{r}(\theta_{l,s},\phi_{l,s},d_{l,s},\mathbf{v})\in\mathbb{C}^{(1\times N_{\text{rx}}\times N_{\text{rz}})}$ is the receiver steering vector, whose formulation is given below
@@ III-A Sensing
The mean square error (MSE) between the true and estimated values is a widely used metric for evaluating sensing performance. However, obtaining a closed-form expression for the MSE is often complex and computationally intensive, as noted in [42]. To address this challenge, we utilise the Cramér-Rao bound (CRB), which provides a theoretical lower bound on the variance of any unbiased estimator. The CRB serves as a benchmark for the minimum achievable MSE, defining the best possible estimation accuracy for a system under ideal conditions.
In the case of an extended target, the BS lacks prior knowledge of the number of scatterers or their respective angles and distances, as echoes are subject to random reflections. Instead of estimating the angles and distances of each individual scatterer, we estimate the entire echo channel matrix $\mathbf{G}$. Accurate estimation of the channel matrix enables the use of algorithms for extracting relevant information. For example, once the channel matrix is accurately estimated, the MUSIC algorithm can be applied to extract the angle information of each scatterer [43, 44].
@@
where $\mathbf{\bar{g}}=\operatorname{vec}\left(\mathbf{G}\right)$ and $\mathbf{\bar{n}}=\operatorname{vec}\left(\mathbf{N}\right)$. According to [45, 46], the FIM of $\mathbf{\bar{g}}$ is given by
@@
In (15), the rank of the matrix $\mathbf{X}$ is $N_{\text{tx}}\times N_{\text{tz}}$, which is sufficient to fully recover the channel matrix $\mathbf{G}$, whose rank is also $N_{\text{tx}}\times N_{\text{tz}}$. As emphasised in [46, 45], a sufficient rank ensures that the FIM remains non-singular. Hence, the CRB of $\mathbf{\bar{g}}$ is given by
@@
Since the FIM of $\mathbf{\bar{g}}$ is independent of the locations of the FAs, the corresponding CRB is also unaffected. However, if the targets are modelled as point targets, as considered in [47, 48], both the FIM and the CRB become dependent on the positions of the FAs. Although a detailed system design of this aspect is beyond the scope of the present work, we provide the analytical derivation to show the dependence of FAs position on CRB for a point target in Appendix B for completeness.
@@ III-B Semantic Communication
$\gamma_{k}$ $=\frac{\mathbb{E}\left[\left|\left|\mathbf{h}_{k}\mathbf{w}_{k}\mathbf{c}^{H}_{k}\right|\right|^{2}\right]}{\sum_{k^{\prime}\in\mathcal{K},k^{\prime}\neq k}\mathbb{E}\left[\left|\left|\mathbf{h}_{k}\mathbf{w}_{k^{\prime}}\mathbf{c}^{H}_{k^{\prime}}\right|\right|^{2}\right]+\mathbb{E}\left[\left|\left|\mathbf{h}_{k}\mathbf{W}_{r}\right|\right|^{2}\right]+\sigma_{c}^{2}}$ (17)
@@
The semantic transmission rate is defined as the number of bits received by the user after decoding the semantic information from the received signal. The expression for the semantic transmission rate is given by [34]:
@@
where the parameter $\rho_{LB}\leq\rho_{k}\leq 1$ represents the semantic extraction ratio shown in (1), and $\iota$ is a scalar value that converts the word-to-bit ratio. Additionally, $\rho_{LB}$ is the lower bound of $\rho_{k}$, with the formula provided in [34, Lemma 1]:
@@
In semantic communication, the privacy of messages is closely linked to the security of the KBs [17]. While shared KBs are essential for enabling efficient semantic encoding and decoding, they also introduce potential privacy vulnerabilities, particularly if an unintended user gains access to a similar or identical KB, which may allow for partial or full reconstruction of the transmitted message [22, 24, 23]. Consequently, ensuring both the confidentiality of the transmitted content and the secure management of KBs is vital. To characterise the associated risks under the worst-case scenario, where KBs are assumed to be exposed (i.e., in the absence of protections such as those mandated by GDPR), we define the worst-case semantic secrecy rate (SSR) as the difference in semantic transmission rates between the legitimate receiver and an unintended user, both assumed to have access to the same KB (i.e., $\rho_{k}=\rho_{l|k}$). Accordingly, the worst-case SSR for the $k$-th CU can be formulated by incorporating (18) and (20):
[…]
@@ III-C Computing
Extracting semantic information from traditional messages largely relies on advanced machine learning techniques, which introduce significant computational overhead. Therefore, it is crucial to consider computational power as a key component of the overall transmission power budget. As outlined in [34], the computational power consumption is modelled using a natural logarithmic function to capture the relationship between computational complexity and power requirements. The formulation is given by:
@@
Moreover, after acquiring the sensing data, analysing and processing these data is crucial for applications such as DTs. This process requires computational resources and power, which must be accounted for. Since the echo signals are received in the same time slot, we consider a parallel computing model, where all computation tasks are computed at the same time. This requires an optimal CPU resource assignment. The computational task associated with the $l$-th target is represented by the tuple $\left(U_{l},Q_{l}\right)$, where $U_{l}$ denotes the size of the sensing data in bits, $Q_{l}$ is the required computational resource in terms of CPU cycles per bit. As demonstrated in various studies, such as [49, 50], the processing latency for such tasks can be expressed as:
@@ IV-A Problem Formulation
The constraints (28h) and (28i) facilitate beamforming design by enforcing single-stream transmission for each user, thereby eliminating the need for advanced signal processing techniques [51]. In the absence of this constraint, more complex transceiver architectures, such as those employing water-filling power allocation or interference cancellation, would be necessary, resulting in significantly higher computational complexity.
In the following section, we relax the rank-one constraint due to its non-convex nature. The rank-one solution can be recovered through Gaussian randomisation [52]. Specifically, under the Gaussian randomisation method, random vectors are generated by sampling from a Gaussian distribution characterised by the obtained high-rank matrix. Each random vector is then evaluated using the original objective function, and the one achieving the best performance is selected as an approximate rank-one solution. To overcome the non-convexity of the objective function in (28), we propose an AO approach. Specifically, we decompose (28) into three sub-problems and iteratively optimise each one until convergence. In the subsequent subsections, we detail each sub-problem and outline the algorithm for solving each sub-problem.
@@ IV-B Joint Beamforming and Computation Optimisation
With given initial FA positions $\mathbf{u}_{i}$ and the initial semantic extraction ratios $\rho_{k}$, we can reformulate the original problem with respect to variables $[\mathbf{w}_{k},\mathbf{R}_{x},f_{l},\zeta]$ as follows:
@@
$\begin{cases}A_{k}=\sum_{k\in\mathcal{K}}\left|\mathbf{h}_{k}\mathbf{w}_{k}\right|^{2}\\ \vspace{0.2cm}\hskip 56.9055pt+\mathbf{h}_{k}\left(\mathbf{R}_{x}-\sum_{k\in\mathcal{K}}\mathbf{w}_{k}\mathbf{w}_{k}^{H}\right)\mathbf{h}_{k}^{H}+\sigma_{c}^{2},\\ B_{k}=\sum_{k^{\prime}\in\mathcal{K},k^{\prime}\neq k}\left|\mathbf{h}_{k}\mathbf{w}_{k^{\prime}}\right|^{2}\\ \vspace{0.2cm}\hskip 56.9055pt+\mathbf{h}_{k}\left(\mathbf{R}_{x}-\sum_{k\in\mathcal{K}}\mathbf{w}_{k}\mathbf{w}_{k}^{H}\right)\mathbf{h}_{k}^{H}+\sigma_{c}^{2},\\ C_{l|k}=\sum_{k^{\prime}\in\mathcal{K},k^{\prime}\neq k}\left|\mathbf{h}_{l}\mathbf{w}_{k^{\prime}}\right|^{2}\\ \vspace{0.2cm}\hskip 56.9055pt+\mathbf{h}_{l}\left(\mathbf{R}_{x}-\sum_{k\in\mathcal{K}}\mathbf{w}_{k}\mathbf{w}_{k}^{H}\right)\mathbf{h}_{l}^{H}+\sigma_{c}^{2},\\ D_{l|k}=\sum_{k\in\mathcal{K}}\left|\mathbf{h}_{l}\mathbf{w}_{k}\right|^{2}\\ \hskip 56.9055pt+\mathbf{h}_{l}\left(\mathbf{R}_{x}-\sum_{k\in\mathcal{K}}\mathbf{w}_{k}\mathbf{w}_{k}^{H}\right)\mathbf{h}_{l}^{H}+\sigma_{c}^{2}.\end{cases}$ (31)
@@
$\begin{cases}\log_{2}\left(B_{k}\right)=\log_{2}\left(B_{k,e}\right)+\frac{1}{B_{k,e}\ln(2)}\left(B_{k}-B_{k,e}\right),\\ \log_{2}\left(D_{l|k}\right)=\log_{2}\left(D_{l|k,e}\right)+\frac{1}{D_{l|k,e}\ln(2)}\left(D_{l|k}-D_{l|k,e}\right),\\ \end{cases}$ (32)
where the subscript $e$ denotes the value of the corresponding variable at each epoch. By substituting (30) and (32), we obtain the expression in (34) to replace (29b). Consequently, the optimisation problem in (29) becomes convex and can be efficiently solved using established optimisation tools such as CVX [53].

### v2

@@ II-A Signal Transmission Model
The semantic communication flow in the proposed framework is shown in Fig. 3. In semantic communication, semantic extraction is primarily guided by the local KBs of the sender and receiver. These local KBs are constructed from a shared KB and enriched with individual knowledge unique to each party. For example, in a vehicular network, the shared KB could be a global map, while each vehicle’s local KB includes this global map along with specific local routes the vehicle has previously traversed. Based on the local KB, the sender focuses on extracting and conveying the underlying meaning of a message. To achieve this, a machine learning-based semantic extraction module is employed to identify key elements that capture the core semantic features of the message. For text-based messages, this typically involves leveraging models such as Transformers to select keywords that maximise performance metrics, such as the similarity between the original message and the reconstructed message at the receiver. The selected words are then encoded using channel coding techniques to ensure robustness against channel noise and transmission errors. A more integrated approach in semantic communication is to employ joint source-channel coding (JSCC), which is a machine learning-based approach such as proposed in [xie2021deep, 10841377], where the transmitted signal $\mathbf{C}$ is generated through a JSCC module $\mathcal{F}(\mathbf{S})$, with $\mathbf{S}$ denoting the source message. Therefore, for each user $k$, the semantic compression ratio is given by
@@
At the receiver’s end, the local KB enables accurate reconstruction of the sender’s intended meaning, even if the exact transmitted words are not received. It is important to note that semantic noise plays a critical role in influencing the performance of semantic inference. Semantic noise may arise from poor channel conditions that cannot be fully mitigated by channel coding and decoding techniques, resulting in semantic distortion. It may also stem from impaired or incomplete KBs, which can lead to misinterpretation of the message [cheng2024knowledge]. To address the former, the proposed framework employs FAs to enhance channel conditions and reduce the impact of transmission-related semantic noise. Semantic noise caused by impaired KBs is beyond the scope of this work, and we assume that the KBs are properly constructed and maintained.
@@
where $\mathbf{W}_{c}\in\mathbb{C}^{(N_{\text{tx}}\times N_{\text{tz}})\times K}$ denote the communication beamforming matrix for the CUs, and each column corresponds to the beamforming vector for an individual user. Specifically, $\mathbf{w}_{k}\in\mathbb{C}^{(N_{\text{tx}}\times N_{\text{tz}})\times 1}$ represents the beamforming vector for the $k$-th user. Thus, the complete beamforming matrix can be expressed as $\mathbf{W}_{c}=[\mathbf{w}_{0},\mathbf{w}_{1},\cdots,\mathbf{w}_{K-1}]$. The semantic communication signals intended for the CUs are denoted by $\mathbf{C}\in\mathbb{C}^{F\times K}$, where each column represents the signal intended for a specific user across $F$ frames. Specifically, $\mathbf{c}_{k}\in\mathbb{C}^{F\times 1}$ denotes the communication signal for the $k$-th CU. Accordingly, the communication signal matrix is given by $\mathbf{C}=[\mathbf{c}_{0},\mathbf{c}_{1},\cdots,\mathbf{c}_{K-1}]$. Additionally, $\mathbf{W}_{r}\in\mathbb{C}^{(N_{\text{tx}}\times N_{\text{tz}})\times F}$ represents the sensing signal. Without loss of generality, if the number of frames $F$ is sufficiently large, we can make the following two assumptions [liuxiang2020joint]:
@@ II-B Communication Model
[…]
@@
[…]
@@
Given that both the CUs and targets are located within the NF region of the BS, the steering vector can be characterised by [liu2023near]:
@@ II-C Sensing Model
[…]
@@ III-A Sensing
The mean square error (MSE) between the true and estimated values is a widely used metric for evaluating sensing performance. However, obtaining a closed-form expression for the MSE is often complex and computationally intensive, as noted in [bekkerman2006target]. To address this challenge, we utilise the Cramér-Rao bound (CRB), which provides a theoretical lower bound on the variance of any unbiased estimator. The CRB serves as a benchmark for the minimum achievable MSE, defining the best possible estimation accuracy for a system under ideal conditions.
In the case of an extended target, the BS lacks prior knowledge of the number of scatterers or their respective angles and distances, as echoes are subject to random reflections. Instead of estimating the angles and distances of each individual scatterer, we estimate the entire echo channel matrix $\mathbf{G}$. Accurate estimation of the channel matrix enables the use of algorithms for extracting relevant information. For example, once the channel matrix is accurately estimated, the MUSIC algorithm can be applied to extract the angle information of each scatterer [li1996adaptive, schmidt1986multiple].
@@
where $\mathbf{\bar{g}}=\operatorname{vec}\left(\mathbf{G}\right)$ and $\mathbf{\bar{n}}=\operatorname{vec}\left(\mathbf{N}\right)$. According to [liu2021cramer, ben2009constrained], the FIM of $\mathbf{\bar{g}}$ is given by
@@
In (15), the rank of the matrix $\mathbf{X}$ is $N_{\text{tx}}\times N_{\text{tz}}$, which is sufficient to fully recover the channel matrix $\mathbf{G}$, whose rank is also $N_{\text{tx}}\times N_{\text{tz}}$. As emphasised in [ben2009constrained, liu2021cramer], a sufficient rank ensures that the FIM remains non-singular. Hence, the CRB of $\mathbf{\bar{g}}$ is given by
@@
Since the FIM of $\mathbf{\bar{g}}$ is independent of the locations of the FAs, the corresponding CRB is also unaffected. However, if the targets are modelled as point targets, as considered in [ma2024movable, chen2025antenna], both the FIM and the CRB become dependent on the positions of the FAs. Although a detailed system design of this aspect is beyond the scope of the present work, we provide the analytical derivation to show the dependence of FAs position on CRB for a point target in Appendix B for completeness.
@@ III-B Semantic Communication
[…]
@@
The semantic transmission rate is defined as the number of bits received by the user after decoding the semantic information from the received signal. The expression for the semantic transmission rate is given by [yang2024secure]:
@@
where the parameter $\rho_{LB}\leq\rho_{k}\leq 1$ represents the semantic extraction ratio shown in (1), and $\iota$ is a scalar value that converts the word-to-bit ratio. Additionally, $\rho_{LB}$ is the lower bound of $\rho_{k}$, with the formula provided in [yang2024secure, Lemma 1]:
@@
In semantic communication, the privacy of messages is closely linked to the security of the KBs [zhaohui2024secure]. While shared KBs are essential for enabling efficient semantic encoding and decoding, they also introduce potential privacy vulnerabilities, particularly if an unintended user gains access to a similar or identical KB, which may allow for partial or full reconstruction of the transmitted message [agarwal2022gdpr, won2024resource, guo2024survey]. Consequently, ensuring both the confidentiality of the transmitted content and the secure management of KBs is vital. To characterise the associated risks under the worst-case scenario, where KBs are assumed to be exposed (i.e., in the absence of protections such as those mandated by GDPR), we define the worst-case semantic secrecy rate (SSR) as the difference in semantic transmission rates between the legitimate receiver and an unintended user, both assumed to have access to the same KB (i.e., $\rho_{k}=\rho_{l|k}$). Accordingly, the worst-case SSR for the $k$-th CU can be formulated by incorporating (18) and (20):
$S_{k}$ $=\max_{l\in\mathcal{L}}\left(R_{k}-R_{l|k},\;0\right)$ (22)
$=\max_{l\in\mathcal{L}}\left(\frac{\iota}{\rho_{k}}\log_{2}\left(\frac{1+\gamma_{k}}{1+\Gamma_{l|k}}\right),\;0\right).$
@@ III-C Computing
Extracting semantic information from traditional messages largely relies on advanced machine learning techniques, which introduce significant computational overhead. Therefore, it is crucial to consider computational power as a key component of the overall transmission power budget. As outlined in [yang2024secure], the computational power consumption is modelled using a natural logarithmic function to capture the relationship between computational complexity and power requirements. The formulation is given by:
@@
Moreover, after acquiring the sensing data, analysing and processing these data is crucial for applications such as DTs. This process requires computational resources and power, which must be accounted for. Since the echo signals are received in the same time slot, we consider a parallel computing model, where all computation tasks are computed at the same time. This requires an optimal CPU resource assignment. The computational task associated with the $l$-th target is represented by the tuple $\left(U_{l},Q_{l}\right)$, where $U_{l}$ denotes the size of the sensing data in bits, $Q_{l}$ is the required computational resource in terms of CPU cycles per bit. As demonstrated in various studies, such as [do2022digital, lu2020communication], the processing latency for such tasks can be expressed as:
@@ IV-A Problem Formulation
The constraints (28h) and (28i) facilitate beamforming design by enforcing single-stream transmission for each user, thereby eliminating the need for advanced signal processing techniques [wu2013physical]. In the absence of this constraint, more complex transceiver architectures, such as those employing water-filling power allocation or interference cancellation, would be necessary, resulting in significantly higher computational complexity.
In the following section, we relax the rank-one constraint due to its non-convex nature. The rank-one solution can be recovered through Gaussian randomisation [wang2014outage]. Specifically, under the Gaussian randomisation method, random vectors are generated by sampling from a Gaussian distribution characterised by the obtained high-rank matrix. Each random vector is then evaluated using the original objective function, and the one achieving the best performance is selected as an approximate rank-one solution. To overcome the non-convexity of the objective function in (28), we propose an AO approach. Specifically, we decompose (28) into three sub-problems and iteratively optimise each one until convergence. In the subsequent subsections, we detail each sub-problem and outline the algorithm for solving each sub-problem.
@@ IV-B Joint Beamforming and Computation Optimisation
With given initial FA positions $\mathbf{u}_{i}$ and the initial semantic extraction ratios $\rho_{k}$, we can reformulate the original problem with respect to variables $\mathbf{w}_{k},\mathbf{R}_{x},f_{l},\zeta]$ as follows:
@@
$\begin{cases}A_{k}=\sum_{k\in\mathcal{K}}\left|\mathbf{h}_{k}\mathbf{w}_{k}\right|^{2}\\ \vskip 5.69046pt\hskip 56.9055pt+\mathbf{h}_{k}\left(\mathbf{R}_{x}-\sum_{k\in\mathcal{K}}\mathbf{w}_{k}\mathbf{w}_{k}^{H}\right)\mathbf{h}_{k}^{H}+\sigma_{c}^{2},\\ B_{k}=\sum_{k^{\prime}\in\mathcal{K},k^{\prime}\neq k}\left|\mathbf{h}_{k}\mathbf{w}_{k^{\prime}}\right|^{2}\\ \vskip 5.69046pt\hskip 56.9055pt+\mathbf{h}_{k}\left(\mathbf{R}_{x}-\sum_{k\in\mathcal{K}}\mathbf{w}_{k}\mathbf{w}_{k}^{H}\right)\mathbf{h}_{k}^{H}+\sigma_{c}^{2},\\ C_{l|k}=\sum_{k^{\prime}\in\mathcal{K},k^{\prime}\neq k}\left|\mathbf{h}_{l}\mathbf{w}_{k^{\prime}}\right|^{2}\\ \vskip 5.69046pt\hskip 56.9055pt+\mathbf{h}_{l}\left(\mathbf{R}_{x}-\sum_{k\in\mathcal{K}}\mathbf{w}_{k}\mathbf{w}_{k}^{H}\right)\mathbf{h}_{l}^{H}+\sigma_{c}^{2},\\ D_{l|k}=\sum_{k\in\mathcal{K}}\left|\mathbf{h}_{l}\mathbf{w}_{k}\right|^{2}\\ \hskip 56.9055pt+\mathbf{h}_{l}\left(\mathbf{R}_{x}-\sum_{k\in\mathcal{K}}\mathbf{w}_{k}\mathbf{w}_{k}^{H}\right)\mathbf{h}_{l}^{H}+\sigma_{c}^{2}.\end{cases}$ (31)
@@
[…]
where the subscript $e$ denotes the value of the corresponding variable at each epoch. By substituting (30) and (32), we obtain the expression in (34) to replace (29b). Consequently, the optimisation problem in (29) becomes convex and can be efficiently solved using established optimisation tools such as CVX [grant2014cvx].